ther are 15 items in a bag. in which 11 are good and 4 are defective . if two items are selected one by one '
1- with replacement.
2- Without replacement'
find the probability that first item is defective and second is good
1
Expert's answer
2020-05-11T18:27:01-0400

1)with replacement

P=(4/15)*(11/(15-1))=0.2095

2)without replacement

P=(4/15)*(11/15)=0.1955